OpenStudy (nuccioreggie):

positive sixteen minus negative thirty-four equals what value negative fifty negative eighteen positive eighteen positive fifty

OpenStudy (nuccioreggie):

@kenzieloo

OpenStudy (nuccioreggie):

@akshay1234

OpenStudy (nuccioreggie):

@SolomonZelman

OpenStudy (hoblos):

it looks like this 16 - (-34)

OpenStudy (paki):

+16-(-34) = 50

OpenStudy (hoblos):

we know that -(-a) = a for any value of a so 16 - (-34) = 16 +34 = ?